### v1

# Revisiting Shooting Point Monte Carlo Methods for Transition Path Sampling

Abstract
Rare event sampling algorithms are essential for understanding processes that occur infrequently on the molecular scale, yet they are important for the long-time dynamics of complex molecular systems. One of these algorithms, transition path sampling, has become a standard technique to study such rare processes since no prior knowledge on the transition region is required. Most TPS methods generate new trajectories from old trajectories by selecting a point along the old trajectory, modifying its momentum in some way, and then “shooting” a new trajectory by integrating forward and backward in time. In some procedures, the shooting point is selected independently for each trial move, but in others, the shooting point evolves from one path to the next so that successive shooting points are related to each other. We provide an extended detailed balance criterion for shooting methods. We affirm detailed balance for most TPS methods, but the new criteria reveals the need for amended acceptance criteria in the flexible length aimless shooting and spring shooting methods.
preprint: APS/123-QED
I Introduction
Transition path sampling (TPS) is a rare events method that yields unbiased insights into the mechanism and kinetics in complex molecular systems, including processes such as nucleation Menzl et al. (2016), biomolecular reorganization ichi Okazaki et al. (2019); Juraszek and Bolhuis (2006) and chemical reactions Basner and Schwartz (2005); Geissler et al. (2001). Unlike many enhanced sampling techniques, TPS does not require detailed a priori knowledge about the process of interest, apart from a definition of the relevant (meta)stable states, i.e., knowledge of the start and end point of the process Bolhuis et al. (2002). The aim of TPS is sampling from a distribution of unbiased dynamical trajectories with start and end points located in these predefined stable states. TPS generates this ensemble of transition paths employing a Markov chain Monte Carlo framework Bolhuis et al. (2002). At the core of many path sampling schemes is the so-called shooting move, in which a new path is generated by the integration of the equations of motion (forward and backward in time) starting from a preselected configuration taken from an existing transition path, known as the shooting point. Valid reactive paths that successfully connect the defined states are accepted/rejected in a Metropolis step. Standard TPS has successfully been used on a wide variety of systems and processes Bolhuis and Dellago (2015); Bolhuis and Swenson (2021).
Several methods have been proposed to increase the sampling efficiency by incorporating prior knowledge of the system into the selection of the shooting point using collective variables, such as biased TPS Bolhuis et al. (2002); Juraszek and Bolhuis (2006) or shooting range TPS Jung et al. (2017). Likewise, artificial intelligence-assisted sampling can increase the sampling efficiency significantly Jung et al. (2023); Falkner et al. (2023) and enable efficient calculation of rates Lazzeri et al. (2023). While methods involving prior knowledge about the reaction constitute a powerful branch of TPS algorithm development, there are still many unknown processes where these might not be directly applicable. For such cases, aimless shooting Peters and Trout (2006); Mullen et al. (2015); Peters (2010) and spring shooting Brotzakis and Bolhuis (2016) were introduced as a system-agnostic, yet efficient, alternatives Beckham and Peters (2011); Joswiak et al. (2018); Arjun and Bolhuis (2021); Janoš et al. (2020). While in most other TPS procedures the shooting point is selected freshly from the current path for each trial move, in the aimless and spring shooting methods, the shooting point location (or index) evolves in a certain way from one path to the next, based on knowledge of the shooting point in the previous step. A central concept in these methods is that this allows the shooting algorithm to focus automatically on the barrier region so that the acceptance is improved.
In this study, we revisit several history-based shooting point Monte Carlo methods, including flexible-length aimless shooting and spring shooting. In these methods, we observe that the particular procedure to select shooting points leads to inconsistencies in the detailed balance equations, resulting in an incorrectly sampled path ensemble. This issue arises because the generation of the new shooting point depends on information on the previous shooting, rendering the procedure non-Markovian. Consequently, MCMC theory does not apply to the trajectory space usually considered in the TPS-formalism.
Therefore, our objective is to develop a general formalism that restores Markovianity and considers the evolving shooting point in the detailed balance equation to ensure correct sampling of the desired path distribution. This formalism makes use of an extended state space that includes shooting points in addition to trajectories. We develop the general formalism and apply it to the case of the aimless and spring shooting algorithms. In particular, we show that the performance of the fixed-length methods depends heavily on the total path length. Furthermore, we find a detailed balance-related issue in the original spring shooting method Brotzakis and Bolhuis (2016), which we are able to correct. One of the consequences of treating the algorithms correctly in the extended space formalism is that they are not behaving as efficiently as concluded in the original papers.
The remaining paper is organized as follows: first, we introduce the extended space of paths and shooting indices, which is required to capture the evolution of the index. We then derive detailed balance relations for Monte Carlo moves in this extended space and apply them to regular TPS, aimless shooting and spring shooting. This leads to corrections for flexible-length aimless and spring shooting that are subsequently tested by performing path sampling in a one-dimensional test system.
II Theory
II.1 Fixed- and Flexible-length Path Ensembles
Assuming two (meta)stable states A and B, we denote the ensemble of paths that connect these states as $P_{\text{AB}}(X)$, where $X=\{x_{1},x_{2},...x_{L}\}$ is a discretized trajectory of length $L$ and each time slice or frame $x$ contains the coordinates (and possibly momenta) of all particles in the system.
The definition of $P_{\text{AB}}(X)$ depends on the desired path ensemble. For instance, a path distribution where all paths are of fixed length $L_{\text{max}}$ can be defined as
$P_{\text{AB}}(X)=\frac{1}{Z_{\text{AB}}}h_{\text{A}}(x_{1})h_{\text{B}}(x_{L})\rho(x_{1})\prod_{i=1}^{L_{\text{max}}-1}p(x_{i}\to x_{i+1})\,,$ (1)
where $h_{\text{A}}(x)$ and $h_{\text{B}}(x)$ are the characteristic functions of the states A and B (unity when $x\in A,B$, zero otherwise), $\rho(x_{1})$ is the stationary distribution generated by the underlying dynamics (e.g. the Boltzmann distribution) and $p(x_{i}\to x_{i+1})$ are the short time transition probabilities. The partition function $Z_{\text{AB}}$ normalizes the distribution.
Another commonly used definition is the flexible-length ensemble, where paths are required to have the start and endpoint in A and B, respectively, while all other points in between should lie outside of A and B:
$P_{\text{AB}}(X)\propto H_{\text{AB}}(X)\prod_{i=1}^{L(X)-1}p(x_{i}\to x_{i+1})\,.$ (2)
Here, $H_{\text{AB}}(X)$ is unity if the path $X$ fulfills all requirements to be considered reactive and zero otherwise, as specified below. Since pathways of different lengths have different measures, one must take special considerations to arrive at a workable distribution. To do so, we consider, in fact, pathways that are of equal length $L_{\text{max}}$ and require that:
  1. 1.
    paths start in A, $x_{1}\in\text{A}$
  2. 2.
    paths reach B, i.e., there is an $i\leq L_{\text{max}}$ such that $x_{i}\in\text{B}$
  3. 3.
    all points of the path except $x_{1}$ are outside of A
If these conditions are met, $H_{\text{AB}}(X)$ is unity and it is zero otherwise. We denote the index of the first point in B as the path length $L(X)$. Note that this way of interpreting the distribution is only formal. In a practical scheme, one halts the integration of the equations of motion when a trajectory enters a stable state. Extending the trajectory (hypothetically) to full length will then result in a valid pathway.
II.2 Extended Space of Paths and Shooting Indices
In path sampling methods like aimless shooting and spring shooting, information on the previous shooting point is used to generate the new shooting point. To take this information into account, we denote with $k$ the index of a (configuration) point $x_{k}$ on the path $X$. This point can either be a shooting point directly or a point that is used to generate the next shooting point depending on the sampling scheme. Independent of the protocol, we refer to $k$ as the shooting index for both cases since, even in the second case, the move is initiated from $k$. Common for both cases is that the index $k$, in some form or the other, evolves from trial to trial. To track the evolution of $X$ and $k$ at the same time in a Monte Carlo procedure and keep the evolution Markovian, we define an extended space consisting of both variables
$Y=(X,k)\,.$ (3)
In each step of the Monte Carlo procedure, a new path $X^{\text{n}}$ is generated together with a new shooting index $k^{\text{n}}$:
$(X^{\text{o}},k^{\text{o}})\to(X^{\text{n}},k^{\text{n}})\,,$ (4)
with a particular generation probability
$p_{\text{gen}}\bigl{[}(X^{\text{o}},k^{\text{o}})\to(X^{\text{n}},k^{\text{n}})\bigr{]}\,,$ (5)
that depends on the details of the generation algorithm. The new path and shooting point are then accepted with acceptance probability $p_{\text{acc}}\bigl{[}(X^{\text{o}},k^{\text{o}})\to(X^{\text{n}},k^{\text{n}})\bigr{]}$. If $(X^{\text{n}},k^{\text{n}})$ is rejected, the old path and shooting point are counted again in the path or index ensembles and the corresponding averages.
In order to derive an acceptance probability for the Monte Carlo move, the joint probability density $P(X,k)$ must be specified. Since our goal is the sampling of trajectories distributed according to $P_{\text{AB}}(X)$, we need to require that marginalizing $P(X,k)$ with respect to $k$ should yield $P_{\text{AB}}(X)$:
$\sum_{k}P(X,k)=P_{\text{AB}}(X)\,.$ (6)
This condition is naturally fulfilled by defining $P(X,k)$ via the conditional probability $p(k|X)$:
$P(X,k)=P_{\text{AB}}(X)p(k|X)\,.$ (7)
Due to the normalization of the conditional probability, $\sum_{k}p(k|X)=1$, the sum over all $k$ is guaranteed to lead to $P_{\text{AB}}(X)$ as marginal distribution. Apart from this condition, there is total freedom in setting the concrete form of $p(k|X)$, the shooting index distribution.
II.3 Shooting Point Monte Carlo
The goal is now to construct a Markov chain Monte Carlo procedure for sampling the joint probability $P(X,k)=P_{\text{AB}}(X)p(k|X)$ from which we can obtain the desired $P_{\text{AB}}(X)$ by marginalizing with respect to $k$. Imposing detailed balance
$P_{\text{AB}}(X^{\text{o}})p(k^{\text{o}}|X^{\text{o}})p_{\text{gen}}\bigl{[}(X^{\text{o}},k^{\text{o}})\to(X^{\text{n}},k^{\text{n}})\bigr{]}$
×$\times p_{\text{acc}}\bigl{[}(X^{\text{o}},k^{\text{o}})\to(X^{\text{n}},k^{\text{n}})\bigr{]}=$
$P_{\text{AB}}(X^{\text{n}})p(k^{\text{n}}|X^{\text{n}})p_{\text{gen}}\bigl{[}(X^{\text{n}},k^{\text{n}})\to(X^{\text{o}},k^{\text{o}})\bigr{]}$
×$\times p_{\text{acc}}\bigl{[}(X^{\text{n}},k^{\text{n}})\to(X^{\text{o}},k^{\text{o}})\bigr{]}$ (8)
yields
$\frac{p_{\text{acc}}\bigl{[}(X^{\text{o}},k^{\text{o}})\to(X^{\text{n}},k^{\text{n}})\bigr{]}}{p_{\text{acc}}\bigl{[}(X^{\text{n}},k^{\text{n}})\to(X^{\text{o}},k^{\text{o}})\bigr{]}}=$
$\frac{P_{\text{AB}}(X^{\text{n}})p(k^{\text{n}}|X^{\text{n}})p_{\text{gen}}\bigl{[}(X^{\text{n}},k^{\text{n}})\to(X^{\text{o}},k^{\text{o}})\bigr{]}}{P_{\text{AB}}(X^{\text{o}})p(k^{\text{o}}|X^{\text{o}})p_{\text{gen}}\bigl{[}(X^{\text{o}},k^{\text{o}})\to(X^{\text{n}},k^{\text{n}})\bigr{]}}\,,$ (9)
which can be satisfied with the Metropolis criterion. The generation probability can be further factorized as
$p_{\text{gen}}\bigl{[}(X^{\text{o}},k^{\text{o}})$ $\to(X^{\text{n}},k^{\text{n}})\bigr{]}=$
$p_{\text{gen}}(X^{\text{n}}|X^{\text{o}},k^{\text{o}})p_{\text{gen}}(k^{\text{n}}|X^{\text{n}},X^{\text{o}},k^{\text{o}})\,,$ (10)
to separate the generation probabilities of the new path from the generation probabilities of the new shooting index. For the derivation here, we assume that all pathways are generated using shooting moves with the same underlying dynamics, which conserve the equilibrium distribution. This leads to a cancellation of the path probabilities $P_{\text{AB}}(X)$ and the respective generation probabilities, where we factorize $p_{\text{gen}}$ as in Eq. II.3 Dellago et al. (2002). For the flexible-length path ensemble this leads to:
$\frac{p_{\text{acc}}\bigl{[}(X^{\text{o}},k^{\text{o}})\to(X^{\text{n}},k^{\text{n}})\bigr{]}}{p_{\text{acc}}\bigl{[}(X^{\text{n}},k^{\text{n}})\to(X^{\text{o}},k^{\text{o}})\bigr{]}}=$
$H_{\text{AB}}(X^{\text{n}})\frac{p(k^{\text{n}}|X^{\text{n}})p_{\text{gen}}(k^{\text{o}}|X^{\text{o}},X^{\text{n}},k^{\text{n}})}{p(k^{\text{o}}|X^{\text{o}})p_{\text{gen}}(k^{\text{n}}|X^{\text{n}},X^{\text{o}},k^{\text{o}})}\,.$ (11)
Below, we will use this equation to derive correct acceptance rules for different TPS-algorithms.
II.4 Shooting Point Densities
To assess the convergence of a path sampling scheme and further understand its inner workings, it is useful to consider the distribution of points on transition paths and the density of points from which shooting moves are initiated. We can derive the density of points on transition paths based on the points $x_{i}$ on paths at indices $i$ and the path density $P_{\text{AB}}(X)$:
[…]
$=\frac{\int\text{d}X\,P_{\text{AB}}(X)\sum_{i=1}^{L(X)}\delta(x-x_{i})}{\int\text{d}X\,P_{\text{AB}}(X)L(X)}$
$=\frac{1}{\langle L(X)\rangle_{X}}\int\text{d}X\,P_{\text{AB}}(X)\sum_{i=1}^{L(X)}\delta(x-x_{i})\,,$ (12)
where the delta function probes the occurrence of $x$ at position $i$ on a transition path $X$. Note that shorter paths contribute fewer points to this distribution. This density is numerically tractable via a histogram of points or collective variables for complex systems and it is a key indicator for the correct convergence of a path sampling scheme.
Figure 1: Monte Carlo moves in the extended space of path $X$ and shooting index $k$. An index shift (left) generates a new shooting index $k^{\text{o}}\to k^{\text{n}}$ without modifying the current path. The shooting move in the extended space generates a new path $X^{\text{o}}\to X^{\text{n}}$ and, in most cases, a new shooting index $k^{\text{o}}\to k^{\text{n}}$. This is due to the fact that indices may be renumbered after the generation of the new path.
Another distribution that is important to discuss is the distribution of attempted shooting points given that one follows the Monte Carlo scheme described in the previous section. In contrast to $\rho(x|\text{TP})$, this distribution depends on the employed protocol. To obtain a general expression for the distribution of attempted shooting points, we need to consider that $k$ does not need to be the shooting index directly but may also be used to generate the next shooting index. For that, we define $p_{\text{sel}}(\hat{k}|X,k)$ as the probability to initiate shooting from another index $\hat{k}$ given the current index $k$, which can be used to define the (protocol-dependent) shooting point distribution:
$\rho(x|\text{SP})=\int\text{d}X\,$ $P_{\text{AB}}(X)\sum_{k=1}^{L(X)}\biggl{[}p(k|X)$
×$\times\sum_{\hat{k}=1}^{L(X)}p_{\text{sel}}(\hat{k}|X,k)\delta(x-x_{\hat{k}})\biggr{]}\,.$ (13)
For path sampling schemes that do not evolve the shooting point from trial to trial, the new shooting point is often generated independently by directly drawing from $p(k|X)$. In that case, the above expression reduces to
$\rho(x|\text{SP})=\int\text{d}X\,P_{\text{AB}}(X)\sum_{k=1}^{L(X)}p(k|X)\delta(x-x_{k})\,.$ (14)
For a uniform shooting index distribution $p(k|X)=\frac{1}{L(X)}$ with independent drawing of shooting indices, as in standard TPS, we arrive at
$\rho(x|\text{SP})=\int\text{d}X\,P_{\text{AB}}(X)\frac{1}{L(X)}\sum_{k=1}^{L(X)}\delta(x-x_{k})\,,$ (15)
which can be transformed to the distribution of points on paths $\rho(x|\text{TP})$ by weighting each path with a factor proportional to $L(X)$.
To gain insights into the relation of the attempted shooting point distributions in equations (II.4), (14) and  (15) to the extended space $(X,k)$, we derive the distribution of points $x_{k}$ located at $k$:
$\rho_{\text{ens}}(x)$ $=\int\text{d}X\,\sum_{k=1}^{L(X)}P(X,k)\delta(x-x_{k})$
$=\int\text{d}X\,\sum_{k=1}^{L(X)}P_{\text{AB}}(X)p(k|X)\delta(x-x_{k})$
$=\int\text{d}X\,P_{\text{AB}}(X)\sum_{k=1}^{L(X)}p(k|X)\delta(x-x_{k})\,.$ (16)
Here, we arrive at an expression equivalent to (14). That means $\rho_{\text{ens}}(x)$, the distribution of points at $k$, is the same as the distribution of attempted shooting points one would obtain if shooting was performed by freshly drawing $k$ each trial from $p(k|X)$. Since $\rho_{\rm ens}(x)$ is independent of the specific shooting scheme yet depends on $p(k|X)$, we use it to assess the convergence of shooting schemes in the extended space $(X,k)$.
II.5 One-way and Two-way TPS
In fixed- or flexible-length regular TPS, one draws the shooting index $k$ freshly at each round. As a result, there is no dependence of the picked candidate shooting index on the previous shooting index $k^{\text{o}}$ and Markovianity is guaranteed without the extended space formalism. Nonetheless, we can construct a move that would be non-Markovian if the shooting index was not tracked yet stays close to the original two-way shooting algorithm. For that, the algorithm can be split into two parts: one move that redraws the shooting index and the shooting move itself (Fig. 1). This way, even though the difference is subtle, a memory of the shooting index is maintained from trial to trial as the shooting move is initiated from the shooting index of the previous trial. As a result, the distribution of shooting points follows Eq. (14). For redrawing the shooting point, we replace the old shooting index $k^{\text{o}}$ with an index $k^{\text{n}}$ independently sampled from the uniform distribution:
$p(k|X)=\frac{1}{L(X)}\ \text{for}\ 1\leq k\leq L\,.$ (17)
This move naturally does not require an acceptance step since it conserves the distribution of $k$ on the old path $X^{\text{o}}$. We then set the drawn shooting index to $k^{\text{o}}$ for the next move. In the second step, a shooting move is initiated that shoots from the point at $k^{\text{o}}$ and produces the path $X^{\text{n}}$. We then denote the index of the common point of $X^{\text{o}}$ and $X^{\text{n}}$ on the new path as $k^{\text{n}}$. Here, the mapping $k^{\text{o}}$ to $k^{\text{n}}$ is unique and reversible with unit generation probability. We derive the acceptance criterion based on Eq. II.3 and Eq. 17 as
$p_{\text{acc}}\bigl{[}(X^{\text{o}},k^{\text{o}})$ $\to(X^{\text{n}},k^{\text{n}})\bigr{]}$
$=H_{\text{AB}}(X^{\text{n}})\min\biggl{[}1,\frac{p(k^{\text{n}}|X^{\text{n}})}{p(k^{\text{o}}|X^{\text{o}})}\biggr{]}$
$=H_{\text{AB}}(X^{\text{n}})\min\biggl{[}1,\frac{L(X^{\text{o}})}{L(X^{\text{n}})}\biggr{]}\,,$ (18)
which is identical to the acceptance criterion derived previously Dellago et al. (2002); Juraszek and Bolhuis (2006).
In the same way, the shooting range algorithm Jung et al. (2017) can be formulated in the extended ensemble. In that case, shooting indices $k$ are only selected in a range defined by a collective variable $r(x)$:
$p(k|X)=\frac{h(r(x_{k}))}{\sum_{i}h(r(x_{i}))}\,,$ (19)
where $h(r)$ is unity if $a\leq r(x)\leq b$ with user defined bounds $a$ and $b$ and zero otherwise. The derived acceptance includes the above shooting index distribution and the condition must hold that all paths have at least one point within the shooting range Jung et al. (2017). Alternatively, a weight function $w$ can be defined for each $k$ to bias the shooting index selection
$p(k|X)=\frac{w(r(x_{k}))}{\sum_{i}w(r(x_{i}))}\,.$ (20)
Indeed, setting $w(r)=h(r)$ recovers eq (19). In this case, the acceptance criterion becomes
$p_{\text{acc}}\bigl{[}(X^{\text{o}},\hat{k}^{\text{o}})$ $\to(X^{\text{n}},\hat{k}^{\text{n}})\bigr{]}$
$=H_{\text{AB}}(X^{\text{n}})\min\biggl{[}1,\frac{\sum_{i}w(r(x^{\text{o}}_{i}))}{\sum_{i}w(r(x^{\text{n}}_{i}))}\biggr{]}\,.$ (21)
II.6 Fixed-length Aimless Shooting
Aimless shooting with fixed path lengths is historically the basis for algorithms that evolve the shooting point from trial to trial. While there are different formulations of the algorithm in literature, we focus here on the two-point variant described in reference Peters (2010). Here, we start with an initial path $X^{\text{o}}$ of fixed path length $L$. The shooting index distribution is only nonzero at two-points, $L/2-\Delta k$ and $L/2+\Delta k$, where $\Delta k$ is an adjustable parameter:
$p(k|X)=\begin{cases}\frac{1}{2}\ \text{if }k\in\{\frac{L}{2}-\Delta k,\frac{L}{2}+\Delta k\},\\ 0\ \text{otherwise}\end{cases}\,.$ (22)
The two-point aimless algorithm starts by selecting one of these two indices on the old path as a shooting index, which one can describe as an index redrawing as above for regular TPS. Then, integration is performed in the backward direction by $L/2-\Delta k$ and in the forward direction by $L/2+\Delta k$ to obtain $X^{\text{n}}$. This produces a new path on which the index of the common point $\hat{k}^{\text{n}}$ can be either at index $L/2-\Delta k$ if the path goes from state A to state B or at index $L/2+\Delta k$ for a state B to state A path. In equilibrium, both of these outcomes have equal probability and therefore the new shooting index follows the desired $p(k|X)$. Due to the splitting of the generation of the new shooting index and the shooting move itself, the acceptance for the shooting move only includes the shooting index distributions
$p_{\text{acc}}\bigl{[}(X^{\text{o}},\hat{k}^{\text{o}})$ $\to(X^{\text{n}},\hat{k}^{\text{n}})\bigr{]}$
$=H_{\text{AB}}(X^{\text{n}})\min\biggl{[}1,\frac{p(\hat{k}^{\text{n}}|X^{\text{n}})}{p(\hat{k}^{\text{o}}|X^{\text{o}})}\biggr{]}$
$=H_{\text{AB}}(X^{\text{n}})\,.$ (23)
It is important to note that the shooting point does not evolve in fixed-length aimless shooting compared to its flexible-length counterpart. While there is a chance that the next shooting move is initiated from the same point as in the last trial, the shooting index is always drawn independently as $L/2-\Delta k$ or $L/2+\Delta k$ each trial, effectively erasing the memory of the previous shooting index. The distribution of the shooting points from which a shoot is attempted is according to Eq. II.4:
$\rho(x|\text{SP})=\int\mathrm{d}X\,P_{\text{AB}}(X)\frac{1}{2}[$ $\delta(x_{L/2-\Delta k}-x)$
$+\delta(x_{L/2+\Delta k}-x)]\,,$ (24)
which is the distribution of points ±$\pm\Delta k$ away from the path centers. We validate this result via numerical simulations in section III.
II.7 Flexible-length Aimless Shooting
Figure 2: The two-point aimless shooting move in the extended space of paths $X$ and shooting indices $k$. (1) The move starts with the current path $X^{\text{o}}$, shooting index $k^{\text{o}}$ and direction $s^{\text{o}}$ that determines the position of the valid shooting indices. (2) The first index shift is performed by selecting one of the two potential shooting indices which generates $\hat{k}^{\text{o}}$. (3) An extended space shooting move generates $X^{\text{n}}$ and $\hat{k}^{\text{n}}$. At the same time, a new direction $s^{\text{n}}$ is drawn that determines if the second potential shooting point is on the forward or backward segment of the new path. (4) A second index shift is performed leading to $k^{\text{n}}$. This shift is required to ensure the reverse move is possible (tracing bottom to top).
Flexible-length aimless shooting, as proposed by Mullen et al. Mullen et al. (2015), generalizes the aimless shooting move to a flexible-length scenario, which is particularly useful in systems with diffusive dynamics. Similar to fixed-length aimless shooting, the shooting index for the following trial is chosen from a narrow set of possible indices. Due to the flexible-length setting, these points are no longer defined around the path center as in Eq. 22, but close to the previous successful shooting point. The proposed acceptance criterion does not include a ratio of path lengths Mullen et al. (2015).In the following, we will show, based on our new formalism, that such a factor is necessary to sample the correct transition path ensemble.
Using the proposed shooting point Monte Carlo formalism, we see a crucial change going from fixed- to flexible-length moves. Namely, the new shooting index $k^{\text{n}}$ is selected depending on the previous index $k^{\text{o}}$ (Fig. 2). In particular, the two-point variant described in Mullen et al. (2015) samples the new shooting index from $\{k^{\text{o}},k^{\text{o}}+s\Delta k\}$, where $s\in\{-1,1\}$ defines whether the second shooting point is on the backward or forward segment. Since the sign $s$ is preserved until a next trial is accepted, we include it in the extended ensemble $(X,k,s)$ (see SI for detailed balance derivation). To respect reversibility, the generation probability of $k^{\text{n}}$ is divided into two steps. We first choose a shooting index $\hat{k}^{\text{o}}$ on the old path according to:
$p_{\text{gen}}(\hat{k}^{\text{o}}|X^{\text{o}},k^{\text{o}},s^{\text{o}})=\begin{cases}\frac{1}{2}\ \text{if }\hat{k}^{\text{o}}\in\{k^{\text{o}},k^{\text{o}}+s^{\text{o}}\Delta k\},\\ 0\ \text{else}\end{cases}\,.$ (25)
The new path $X^{\text{n}}$ is generated by shooting from $\hat{k}^{\text{o}}$ and integrating the equations of motion until a stable state is reached. We then obtain a new sign $s^{\text{n}}$ by drawing from $\{-1,1\}$ with equal probability. We assume $p(s^{\text{n}})=p_{\text{gen}}(s^{\text{n}})=\frac{1}{2}$. Denoting $\hat{k}^{\text{n}}$ as the index of the shooting point on $X^{\text{n}}$, the new shooting index $k^{\text{n}}$ is generated according to:
$p_{\text{gen}}(k^{\text{n}}|$ $X^{\text{n}},\hat{k}^{\text{n}},s^{\text{n}},X^{\text{o}},k^{\text{o}})=$
$\begin{cases}\frac{1}{2}\ \text{if }k^{\text{n}}\in\{\hat{k}^{\text{n}},\hat{k}^{\text{n}}+s^{\text{n}}\Delta k\},\\ 0\ \text{else}\end{cases}\,.$ (26)
This two-step procedure ensures that the path $X^{\text{o}}$ and shooting index $k^{\text{o}}$ can be generated from $\{X^{\text{n}}$, $k^{\text{n}}\}$ in the reverse move. To simplify the notation, we summarize the $1/2$ generation probabilities from Eqs. (25), (II.7) and $p_{\text{gen}}(s)$ as:
$p_{\text{gen}}(k^{\text{n}}|X^{\text{n}},X^{\text{o}},k^{\text{o}},s^{\text{n}})\,p_{\text{gen}}(s^{\text{n}})$
$=p_{\text{gen}}(k^{\text{o}}|X^{\text{o}},X^{\text{n}},k^{\text{n}},s^{\text{o}})\,p_{\text{gen}}(s^{\text{o}})=\frac{1}{8}\,.$ (27)
When inserting these generation probabilities into the analogous form of Eq. II.3 derived in the SI, we obtain
$\frac{p_{\text{acc}}\bigl{[}(X^{\text{o}},k^{\text{o}})\to(X^{\text{n}},k^{\text{n}})\bigr{]}}{p_{\text{acc}}\bigl{[}(X^{\text{n}},k^{\text{n}})\to(X^{\text{o}},k^{\text{o}})\bigr{]}}=H_{\text{AB}}(X^{\text{n}})\frac{p(k^{\text{n}}|X^{\text{n}})}{p(k^{\text{o}}|X^{\text{o}})}\,.$ (28)
Hence, we are left with the ratio of the shooting index distributions, which are not included in the original algorithm Mullen et al. (2015). As a result, accepting a path if $H_{\text{AB}}(X^{\text{n}})=1$ in a flexible-length setting violates detailed balance, which is also apparent in the results of numerical simulations (see section III). However, following the described two-step procedure, a properly weighted path ensemble can be obtained even a posteriori by assigning each path a statistical weight of $\Omega(X)=1/L(X)$ under the assumption of a uniform shooting index distribution (see SI for derivation).
II.8 Spring Shooting
Inspired by the flexible-length aimless shooting scheme, the work by Brotzakis et al. Brotzakis and Bolhuis (2016) aimed to achieve a similar efficiency gain in a one-way shooting framework. The displacement $\Delta k$ from the previous shooting index is chosen according to an exponential distribution, i.e., a Boltzmann distribution based on a linear potential (constant force), hence the scheme is named spring shooting. We reformulate the algorithm in the shooting point MC framework (Fig. 3), starting out by selecting the direction of the shooting move and setting $s=-1$ for a forward and $s=1$ for a backward shoot. Note that in this case, $s$ is redrawn at the start of every trial and therefore does not have to be included in the extended space. The displacement $\Delta k$ is then sampled according to:
$\pi(\Delta k)=\frac{\min[1,\exp(s\sigma\Delta k)]}{\sum_{\tau=-\Delta k_{\text{max}}}^{\Delta k_{\text{max}}}\min\bigl{[}1,\exp(s\sigma\tau)\bigr{]}}\,,$ (29)
where $\sigma$ is the strength of the spring and $\Delta k_{\text{max}}$ is the maximum displacement. Following the notation of flexible-length aimless shooting, the shooting index $\hat{k}^{\text{o}}$ is set as $k^{\text{o}}+\Delta k$. The generation probability of the new shooting index becomes
$p_{\text{gen}}(\hat{k}^{\text{o}}|X^{\text{o}},k^{\text{o}})=\begin{cases}\pi(\hat{k}^{\text{o}}-k^{\text{o}})\ \text{if }|\hat{k}^{\text{o}}-k^{\text{o}}|\leq\Delta k_{\text{max}},\\ 0\ \text{else}\end{cases}\,.$ (30)
Then, the new path $X^{\text{n}}$ is generated from the configuration at $\hat{k}^{\text{o}}$. The index of the common point of $X^{\text{n}}$ and $X^{\text{o}}$ on the new path is denoted as $\hat{k}^{\text{n}}$. As for aimless shooting, we then shift the new shooting index $\hat{k}^{\text{n}}$ a second time by sampling a shift $\Delta k$ from
$\tilde{\pi}(\Delta k)=\frac{\min[1,\exp(-s\sigma\Delta k)]}{\sum_{\tau=-\Delta k_{\text{max}}}^{\Delta k_{\text{max}}}\min\bigl{[}1,\exp(-s\sigma\tau)\bigr{]}}\,,$ (31)
which corresponds to the following generation probability:
$p_{\text{gen}}(k^{\text{n}}|X^{\text{n}},\hat{k}^{\text{n}})=\begin{cases}\tilde{\pi}(k^{\text{n}}-\hat{k}^{\text{n}})\ \text{if }|k^{\text{n}}-\hat{k}^{\text{n}}|\leq\Delta k_{\text{max}},\\ 0\ \text{else}\end{cases}\,.$ (32)
This means that, given a forward shot, first the shooting index is shifted preferably backwards and, after generating the new path, the new shooting index is generated by shifting preferably forward. Thereby, the move is reversible and generation probabilities are symmetric (see SI). In this case, the resulting acceptance ratio is identical to Eq. 28 and a shooting index distribution must be defined.
Figure 3: The spring shooting move in the extended space of paths $X$ and shooting indices $k$. (1) The move starts with the current path $X^{\text{o}}$ and shooting index $k^{\text{o}}$. (2) The first index shift generating $\hat{k}^{\text{o}}$ is performed by a weighted selection from neighboring points (illustrated using color gradient). (3) An extended space shooting move in the forward direction generates $X^{\text{n}}$ and $\hat{k}^{\text{n}}$. (4) A second index shift is performed leading to $k^{\text{n}}$. This shift is performed with opposite weights (see Eq. 31) and is required to ensure the reverse move is possible (tracing bottom to top).
In the original algorithm proposed by Brotzakis et al. Brotzakis and Bolhuis (2016), the new shooting index $k^{\text{n}}$ is set implicitly to $\hat{k}^{\text{n}}$, the common point of the new an old path, i.e., the second shifting is omitted. While this formulation of spring shooting allows the reverse move in path space, i.e., regenerating the old path $X^{\text{o}}$ from the new path $X^{\text{n}}$, it does not regenerate the old shooting index $k^{\text{o}}$ from $k^{\text{n}}$. As a result, detailed balance is violated in the extended space framework. As for flexible-length aimless shooting, the ratio of shooting index distributions is not considered in the original form of the sampling scheme. In the case of spring shooting, a reweighting factor for paths sampled without a defined shooting index distribution cannot be derived due to the irreversibility of the shooting move in the extended space.
III Results and Discussion
Figure 4: Fixed-length transition path sampling using aimless shooting. (A) The potential energy function of the model system and corresponding state definitions. (B) The density of points on transition paths for different fixed path lengths. Dashed lines correspond to the reference from a long equilibrium simulation, solid lines to the density obtained via aimless shooting. (C) Shooting point distributions for aimless shooting (solid line) and the points at $L/2\pm\Delta k$ on transition paths from the equilibrium simulation (dashed lines).
Figure 5: Flexible-length transition path sampling using aimless shooting as proposed by Mullen et al. Mullen et al. (2015) and spring shooting as proposed by Brotzakis et al. Brotzakis and Bolhuis (2016) (top row) and with the proposed corrections (bottom row). (A, D) Comparison of the density of points on paths between the equilibrium reference (dashed line) and path sampling algorithms. (B,E) Distribution of points at $k$ in the $(X,k)$ extended ensemble for the different path sampling algorithms. The reference distribution was obtained via reweighting the reference density in panel A with a factor proportional to $1/L(X)$. (C, F) The distribution of path lengths obtained using the different path sampling schemes.
As a reference for both fixed- and flexible-length path ensembles, we perform long equilibrium simulations from which we extract the transition path ensemble from spontaneous barrier crossings. All simulations are run employing overdamped Langevin dynamics in a one-dimensional, asymmetric double-well model (Fig. 4A) with a potential energy defined as (all simulation parameters in SI):
$U(x)=\begin{cases}0.2(x-1)^{2}\bigl{[}0.01(x-1)^{2}-1\bigr{]}\ \text{if }x<1,\\ 0.2(x-1)^{2}\bigl{[}0.16(x-1)^{2}-4\bigr{]}\ \text{else}\end{cases}\,.$ (33)
This barrier is about $5\,k_{\text{B}}T$ high. We chose an asymmetric barrier rather than a symmetric one to highlight differences in the obtained shooting point distributions.
For the fixed-length path ensemble (Fig. 4B,C), we compare the density of points on paths $\rho(x|\text{TP})$ as in Eq. II.4 with the shooting point distribution as defined in Eq. II.4. The density of points in the transition region between states A and B decreases for longer paths as less time is spent in the barrier region. This in turn leads to the points at $L/2\pm\Delta k$ being less constrained to be on the barrier top as the time at which the path leaves state A becomes increasingly variable (Fig. 4C). For the longest paths in these simulations, there is hardly any difference between the density of points at $L/2\pm\Delta k$ and $\rho(x|\text{TP})$. Since aimless shooting in a fixed-length setting samples these points as shooting points, the performance of the shooting scheme will depend on the chosen path length. For long paths, both uniform, two-way and aimless shooting initiate shooting moves effectively from the same shooting point distribution. As the shooting index is independently drawn at each trial and the shooting point distribution follows Eq. II.6, the efficiency of fixed-length aimless shooting originates from shooting close to path centers rather than from a restoring force on the shooting point that directs it to the barrier top. For long path lengths, the acceptance of the shooting move suffers as shooting points are primarily taken from the stable states rather than the barrier region.
Continuing with the flexible-length path ensemble, we see that uniform two-way shooting converges to the reference ensemble, indicated by overlapping $\rho(x|\text{TP})$ distributions and path length distributions $p(L)$ (Fig. 5). The density of points at $k$ noted as $\rho_{\text{ens}}(x)$ matches the reweighted density of points on paths (Eq. 15) obtained from the reference trajectory since shooting indices are selected with a uniform $p(k|X)$ on each path. In contrast, flexible-length aimless shooting as proposed by Mullen et al. Mullen et al. (2015) and spring shooting as proposed by Brotzakis et al. Brotzakis and Bolhuis (2016) do not converge to the reference, as can be seen in panels A and C of Fig. 5. Panel B shows that the shooting points for spring shooting are much more focused around the location of the barrier. While this is a desired feature, it clearly gives rise to the wrong distribution of transition paths (panels A and C). We note that the deviations from the reference path ensemble are less clearly seen at higher barriers, where the distribution of shooting points is more localized for all methods. This is also the reason why the spring shooting led to seemingly converged results in Brotzakis and Bolhuis (2016).
Including the shooting index distribution $p(k|X)$ in the acceptance and, for spring shooting, performing a second shift of the shooting index, recovers the convergence to the reference path ensemble (panels D, E, and F of Fig. 5). However, since we impose a uniform $p(k|X)$, shooting moves are now initiated from the same points as in standard uniform two-way shooting. In addition, the small displacements of the shooting index in aimless and spring shooting introduce correlations between shooting points that are absent in uniform one- or two-way shooting. These correlations lead to a slow exploration of shooting points away from the barrier top. For sufficiently large barriers, this can lead to disproportionate shooting from the barrier top in the early stages of sampling, which in turn leads to a slow convergence to the correct path ensemble.
IV Conclusion
We propose a theoretical framework to describe transition path sampling in an extended ensemble including paths and respective shooting indices. Revisiting algorithms that evolve the shooting point from trial to trial, we derive detailed balance equations based on the extended ensemble leading to corrections for aimless and spring shooting in a flexible-length setting. Although the deviation of path ensembles sampled via aimless shooting or spring shooting from the correct path ensemble is system-dependent, a significant overlap in the distributions was found to remain in our experiments. This indicates that previous works may be critically assessed and qualitative conclusions likely still hold. With recent advances in interface-based sampling schemes such as stone skipping Riccardi et al. (2017) and wire fencing Zhang et al. (2023) for transition interface sampling and virtual interface exchange Brotzakis and Bolhuis (2019) for transition path sampling, we hope to see further development of shooting schemes for TPS that require minimal information about the system of interest.
Acknowledgements.
This research was funded in part by the Austrian Science Fund (FWF) 10.55776/F81. For Open Access purposes, the author has applied a CC BY public copyright license to any author accepted manuscript version arising from this submission.
Supplementary Information: Revisiting Shooting Point Monte Carlo Methods for Transition Path Sampling
(Dated: August 6, 2024)
Simulation Details
All simulations are performed in a one-dimensional model system with the potential energy $U(x)$ defined as:
$U(x)=\begin{cases}0.2(x-1)^{2}\bigl{[}0.01(x-1)^{2}-1\bigr{]}\ \text{if }x<1,\\ 0.2(x-1)^{2}\bigl{[}0.16(x-1)^{2}-4\bigr{]}\ \text{else}\end{cases}\,.$ (S1)
We impose overdamped langevin dynamics with $k_{\text{B}}T=1$ and propagate the system using
$x_{i+1}=x_{i}+\Delta tDf(x)+\sqrt{2\Delta tD}\,\zeta\,,$ (S2)
where $f(x)$ is the force at point $x$ and $\zeta$ is a random variable drawn from a standard normal distribution. For all simulations, we set $\Delta tD=0.01$. For transition path sampling, we assume a point to be in state A for $x<-5$ and in state B for $x>4$. We accumulate $24$ independent simulations each with $5\times 10^{5}$ path sampling trials of which $10^{3}$ are discarded as equilibration period. Aimless and spring shooting are always performed with a maximum displacement of $\Delta k_{\text{max}}=25$ and, for spring shooting, the spring strength is set to $\sigma=0.1$.
Extended Ensemble with additional Variables
We can derive, analogous to the derivation of the Monte Carlo procedure for sampling $(X,k)$, a scheme for sampling the ensemble $(X,k,s)$, where $s$ may describe a sign for selecting a shooting index or even a direction to integrate the equations of motion (as for one-way shooting). First, we define an extended space as:
$Y=(X,k,s)\,.$ (S3)
In each step, a new path $X^{\text{n}}$, shooting point $k^{\text{n}}$ and direction $s$ is generated:
$(X^{\text{o}},k^{\text{o}},s^{\text{o}})\to(X^{\text{n}},k^{\text{n}},s^{\text{n}})\,.$ (S4)
with a generation probability
$p_{\text{gen}}\bigl{[}(X^{\text{o}},k^{\text{o}},s^{\text{o}})\to(X^{\text{n}},k^{\text{n}},s^{\text{n}})\bigr{]}\,.$ (S5)
As for the $(X,k)$ ensemble, we define $P(X,k,s)$ in such a way that marginalizing with respect to $k$ and $s$ should yield $P_{\text{AB}}(X)$:
$\sum_{k}\sum_{s}P(X,k,s)=P_{\text{AB}}(X)\,.$ (S6)
This condition is fulfilled by:
$P(X,k,s)=P_{\text{AB}}(X)p(k|X)p(s)\,,$ (S7)
where a sum over all $k$ and $s$ is guaranteed to lead to $P_{\text{AB}}(X)$ as marginal distribution. Detailed balance for a Markov chain Monte Carlo procedure that samples the joint probability $P(X,k,s)=P_{\text{AB}}(X)p(k|X)p(s)$ is given by:
$P_{\text{AB}}(X^{\text{o}})p(k^{\text{o}}|X^{\text{o}})p(s^{\text{o}})p_{\text{gen}}\bigl{[}(X^{\text{o}},k^{\text{o}},s^{\text{o}})\to(X^{\text{n}},k^{\text{n}},s^{\text{n}})\bigr{]}p_{\text{acc}}\bigl{[}(X^{\text{o}},k^{\text{o}},s^{\text{o}})\to(X^{\text{n}},k^{\text{n}},s^{\text{n}})\bigr{]}$ (S8)
$=P_{\text{AB}}(X^{\text{n}})p(k^{\text{n}}|X^{\text{n}})p(s^{\text{n}})p_{\text{gen}}\bigl{[}(X^{\text{n}},k^{\text{n}},s^{\text{n}})\to(X^{\text{o}},k^{\text{o}},s^{\text{o}})\bigr{]}p_{\text{acc}}\bigl{[}(X^{\text{n}},k^{\text{n}},s^{\text{n}})\to(X^{\text{o}},k^{\text{o}},s^{\text{o}})\bigr{]}\,,$ (S9)
yields
$\frac{p_{\text{acc}}\bigl{[}(X^{\text{o}},k^{\text{o}},s^{\text{o}})\to(X^{\text{n}},k^{\text{n}},s^{\text{n}})\bigr{]}}{p_{\text{acc}}\bigl{[}(X^{\text{n}},k^{\text{n}},s^{\text{n}})\to(X^{\text{o}},k^{\text{o}},s^{\text{o}})\bigr{]}}=\frac{P_{\text{AB}}(X^{\text{n}})p(k^{\text{n}}|X^{\text{n}})p(s^{\text{n}})p_{\text{gen}}\bigl{[}(X^{\text{n}},k^{\text{n}},s^{\text{n}})\to(X^{\text{o}},k^{\text{o}},s^{\text{o}})\bigr{]}}{P_{\text{AB}}(X^{\text{o}})p(k^{\text{o}}|X^{\text{o}})p(s^{\text{o}})p_{\text{gen}}\bigl{[}(X^{\text{o}},k^{\text{o}},s^{\text{o}})\to(X^{\text{n}},k^{\text{n}},s^{\text{n}})\bigr{]}}\,.$ (S10)
Reweighting Aimless Path Ensembles
In the Main Text, we showed that the path ensemble resulting from flexible length aimless shooting as proposed in Mullen et al. (2015) is not properly weighted. This is best illustrated in Figure 2a, where both the spring and aimless algorithms do not recover the reference $p(x|TP)$ distributions.
Here, we show that this improper sampling can be easily corrected by introducing a simple reweighting factor for each of the generated paths in the ensemble. The proper detailed balance equation with symmetric generation probabilities is Eq. (28) in the Main Text
$\frac{p_{\text{acc}}\bigl{[}(X^{\text{o}},k^{\text{o}})\to(X^{\text{n}},k^{\text{n}})\bigr{]}}{p_{\text{acc}}\bigl{[}(X^{\text{n}},k^{\text{n}})\to(X^{\text{o}},k^{\text{o}})\bigr{]}}=H_{\text{AB}}(X^{\text{n}})\frac{p(k^{\text{n}}|X^{\text{n}})}{p(k^{\text{o}}|X^{\text{o}})}\,.$ (S11)
However, the aimless shooting algorithm as proposed in Mullen et al. (2015) uses a different acceptance
$\frac{p_{\text{acc}}\bigl{[}(X^{\text{o}},k^{\text{o}})\to(X^{\text{n}},k^{\text{n}})\bigr{]}}{p_{\text{acc}}\bigl{[}(X^{\text{n}},k^{\text{n}})\to(X^{\text{o}},k^{\text{o}})\bigr{]}}=H_{\text{AB}}(X^{\text{n}})\,.$ (S12)
Clearly, this would, in general, lead to different path statistics. The simplest way to correct for this is to introduce a constant weighting factor $\Omega(X)$ for each sampled path. This would mean that we effectively correct the aimless acceptance criterion as if we had used the following acceptance rule
$\frac{p_{\text{acc}}\bigl{[}(X^{\text{o}},k^{\text{o}})\to(X^{\text{n}},k^{\text{n}})\bigr{]}}{p_{\text{acc}}\bigl{[}(X^{\text{n}},k^{\text{n}})\to(X^{\text{o}},k^{\text{o}})\bigr{]}}=H_{\text{AB}}(X^{\text{n}})\frac{\Omega(X^{\rm n})}{\Omega(X^{\rm o})}\,.$ (S13)
The simplest way to ensure that this is identical to the proper acceptance rule in Eq. (S11) is to set the shooting point distribution $p(k|X)$ to
$p(k|X)=\Omega(X)\,.$ (S14)
Since this should be true for all values of k, we require that the marginal of $P(X,k)$ in the reweighted flexible aimless path ensemble will be equal to the marginal of $P(X,k)$ in the correct ensemble That is,
$P_{\text{AB}}(X)=\sum_{k}P_{\text{AB}}(X)p(k|X)=\sum_{k}P_{\text{AB}}(X)\Omega(X)=P_{\text{AB}}(X)\Omega(X)L(X)=P_{\text{AB}}(X)\,.$ (S15)
This can only be achieved by assigning each path the statistical weight
$\Omega(X)=\frac{1}{L(X)}\,.$ (S16)
So, by reweighting each sampled path with a factor equal to the inverse of the length of the path, the bias in the path ensemble can be corrected. This is only possible when the two-step shifting of the shooting point is followed rigorously, and therefore, no reweighting factor can be derived for the original spring shooting as proposed in SBrotzakis and Bolhuis (2016). Still, when we use the corrected spring method described in the next section, and use the $1/L(X)$ reweighting factor, we do obtain the correct ensemble, as shown in the Main Text.
The above derivation starts with the premise that the flexible aimless algorithm samples the wrong distribution. However, the degree to which the distribution deviates depends on the maximum displacement $\Delta k_{\text{max}}$. We noticed that for the flexible-length aimless algorithm using a range of $\Delta k\in[-\Delta k_{\text{max}},\Delta k_{\text{max}}]\}$ the distribution of points on paths converges to the reference ensemble for large $\Delta k_{\text{max}}$, while it is still off for small $\Delta k_{\text{max}}$.
To understand this, we construct the following argument. When doing a (small) shift in shooting point position as one would do in aimless shooting, one will (almost) always remain on the path itself (that is, the new shooting point lies on the path $X$), and one proceeds with the shooting move. However, when $\Delta k_{\text{max}}$ becomes sufficiently large so that the new trial shooting point index falls outside of the path (of length $L(X)$), and is effectively in one of the stable states, the entire move has to be rejected. In that case one has to recount the current path in the Markov chain.
Assuming $\Delta k_{\text{max}}$ larger than the path length $L(X)$, we can argue that one is effectively choosing a random point on a stretch of the path of length $\Delta_{p}=2\Delta k_{\text{max}}+1$. If $\Delta k_{\text{max}}$ is sufficiently large, the estimate for the chance of falling outside the (flexible) path for this random move in shooting point index is the $(\Delta_{p}-L(X))/\Delta_{p}=1-L(X)/\Delta_{p}$. As one has to reject the move, the path is recounted and the shooting point selection is repeated until one hits a trial shooting point within the length $L(X)$ region. The probability to reject $n$ times in a row is $(1-L(X)/\Delta_{p})^{n}$.
The expected number of rejected moves can be computed as follows: Denoting $x\equiv 1-L(X)/\Delta_{p}$ the expected number of rejected moves is
$\langle n\rangle=\frac{\sum_{n=1}^{\infty}nx^{n}}{\sum_{n=1}^{\infty}x^{n}}\,.$ (S17)
The geometrical sums give
$\langle n\rangle=\frac{\frac{1}{(1-x)^{2}}}{\frac{1}{1-x}}=\frac{1}{1-x}=\frac{1}{L(X)/\Delta_{p}}=\frac{\Delta_{p}}{L(X)}\,.$ (S18)
So in a proper sampling scheme, on average one has to recount the old path in the Markov chain $\Delta_{p}/L(X)$ times. As the $\Delta_{p}$ is just a constant, this gives a relative weighting factor of $1/L(X)$, as expected. When one only has a small $\Delta_{p}$ or $\Delta k_{\text{max}}$, the rejection step almost never happens and all paths get the same weight. This explains why for small (maximum) displacement in the shooting point index the weighting of the paths are not taking into account correctly, while for larger displacement the paths ensemble converge to the correct distribution.
Symmetric Generation Probabilities in Spring Shooting
In the revised spring shooting algorithm, we first draw a displacement from the distribution:
$\pi(\Delta k)=\frac{\min[1,\exp(s\sigma\Delta k)]}{\sum_{\tau=-\Delta k_{\text{max}}}^{\Delta k_{\text{max}}}\min\bigl{[}1,\exp(s\sigma\tau)\bigr{]}}\,.$ (S19)
After generating the new path, we set $\hat{k}^{\text{n}}$ to the index of the shooting point on the new path. This index is then shifted a second time with a displacement given by:
$\tilde{\pi}(\Delta k)=\frac{\min[1,\exp(-s\sigma\Delta k)]}{\sum_{\tau=-\Delta k_{\text{max}}}^{\Delta k_{\text{max}}}\min\bigl{[}1,\exp(-s\sigma\tau)\bigr{]}}\,.$ (S20)
For given $\Delta k_{1}$ from $\pi(\Delta k)$ and $\Delta k_{2}$ from $\tilde{\pi}(\Delta k)$, the generation probability for the whole move is:
$p_{\text{gen}}(k^{\text{n}}|X^{\text{n}},X^{\text{o}},k^{\text{o}})=\pi(\Delta k_{1})\tilde{\pi}(\Delta k_{2})\,.$ (S21)
The generation probability of the reverse move is then given by:
$p_{\text{gen}}(k^{\text{o}}|X^{\text{o}},X^{\text{n}},k^{\text{n}})=\pi(-\Delta k_{2})\tilde{\pi}(-\Delta k_{1})\,,$ (S22)
and the ratio of generation probabilities is then:
$\frac{p_{\text{gen}}(k^{\text{n}}|X^{\text{n}},X^{\text{o}},k^{\text{o}})}{p_{\text{gen}}(k^{\text{o}}|X^{\text{o}},X^{\text{n}},k^{\text{n}})}=\frac{\pi(\Delta k_{1})\tilde{\pi}(\Delta k_{2})}{\pi(-\Delta k_{2})\tilde{\pi}(-\Delta k_{1})}\,.$ (S23)
Since $\pi$ and $\tilde{\pi}$ relate via
$\pi(\Delta k)=\tilde{\pi}(-\Delta k)\,,$ (S24)
the generation probabilities are symmetric
$\frac{p_{\text{gen}}(k^{\text{n}}|X^{\text{n}},X^{\text{o}},k^{\text{o}})}{p_{\text{gen}}(k^{\text{o}}|X^{\text{o}},X^{\text{n}},k^{\text{n}})}=1\,.$ (S25)
Transition Path Sampling Algorithms
Here, we provide pseudo-code for the TPS algorithms discussed in the main text.
List of algorithms I Two-Way TPS (left) and one-Way TPS (right). The function IntegrateToState generates a trajectory from the given point until a stable state is hit. The forward and reverse segments are fused by ConcatenatePath.
List of algorithms II Aimless shooting (left) and spring shooting (right) including the corrections from the extended space formalism. The function IntegrateToState generates a trajectory from the given point until a stable state is hit. The forward and reverse segments are fused by ConcatenatePath. When a displacement is added to a shooting point index, we check if $1\leq k\leq L(X)$ using InBounds, otherwise the move is rejected. For spring shooting, SampleDisplacement draws a $\Delta k$ according to a weight given by equation 29 in the main text.